# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ 3.1 General case
Let $\mathbb{k}$ be a field. For any augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ to $\mathbb{k}$,
$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
@@
Since $HC_{0}(K\#3_{1})=\mathcal{A}^{\mathrm{knot}}_{0}(K\#3_{1})/\partial_{K\#3_{1}}(\mathcal{A}^{\mathrm{knot}}_{1}(K\#3_{1}))$, any augmentation $\varepsilon$ to $\mathbb{k}$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ induces a ring homomorphism
$\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(K\#3_{1})\to\mathbb{k}$
such that $\bar{\varepsilon}(r)=\varepsilon(r)$ for every $r\in R$. Composing with $J$ in Lemma 3.6, we obtain a map
$\bar{\varepsilon}\circ J\colon R[X]/(F)\to\mathbb{k}.$
Let us denote $\zeta=\bar{\varepsilon}\circ J(X)$. Since $\bar{\varepsilon}\circ J(F)=0$ and $J$ is a unital $R$-algebra map, we obtain an equation in $\mathbb{k}$
$\zeta^{2}+(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))\zeta+(\varepsilon(\mu)-1)=0.$
Then, $D_{\varepsilon}$ in the lemma is the discriminant of this quadratic equation for $\zeta$. In order that this equation has a solution in $\mathbb{k}$, it is necessary that $D_{\varepsilon}$ is the square of some element of $\mathbb{k}$. ∎
Combining Theorem 3.2, Lemma 3.9 and a computation of the knot DGA of the unknot, we obtain a constraint on the isotopy class of clean Lagrangian intersections.
@@
Let $K_{0}$ be the connected sum of an arbitrary knot $K$ in $\mathbb{R}^{3}$ with $3_{1}$ or its mirror $\overline{3_{1}}$. Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
@@
Assume that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. In the case of $K_{0}=K\#3_{1}$, we can deduce a contradiction as follows: From the trivial braid on a single strand, we can compute as in [16, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
$U-\lambda-\mu+\lambda\mu.$
Let us choose an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{Q}$ determined by
$\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1.\end{array}$
Composing with the map $\Psi_{\varphi}$ in Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{Q}$. It is an augmentation of the knot DGA of $K_{0}=K\#3_{1}$ to $\mathbb{Q}$ such that
$\begin{array}[]{ccc}\varepsilon(\lambda)=3^{a}2^{b}(-1)^{c},&\varepsilon(\mu)=2^{a}(-1)^{d},&\varepsilon(U)=-1.\end{array}$
We apply Lemma 3.9 for $\mathbb{k}=\mathbb{Q}$. Then,
$D_{\varepsilon}=(\varepsilon(\mu)+1+\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)$
must be the square of some rational number. However, we can calculate as follows:
$D_{\varepsilon}=\begin{cases}(\frac{7}{2})^{2}-4\cdot 1=\frac{33}{4}&\text{ if}\varepsilon(\mu)=2,\\ (-\frac{3}{2})^{2}-4\cdot(-3)=\frac{57}{4}&\text{ if }\varepsilon(\mu)=-2,\\ (\frac{7}{2})^{2}-4\cdot(-\frac{1}{2})=\frac{57}{4}&\text{ if }\varepsilon(\mu)=\frac{1}{2},\\ (-\frac{3}{2})^{2}-4\cdot(-\frac{3}{2})=\frac{33}{4}&\text{ if }\varepsilon(\mu)=-\frac{1}{2},\end{cases}$
and this is a contradiction.
In the case of $K_{0}=K\#\overline{3_{1}}$, we can deduce a contradiction as follows: Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
@@
$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ and $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ intersect cleanly along $\overline{K_{1}}$. Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $3_{1}$, this contradicts to the result in the previous case. ∎
@@
For the proof of the theorem, we may choose other field $\mathbb{k}$ and other augmentation $\varepsilon_{1}$ of the knot DGA of $K\#3_{1}$ to $\mathbb{k}$ such that $\varepsilon(U)\neq 1$. Let us see two examples.
  1. 1.
    Choose $\mathbb{k}=\mathbb{Q}[\sqrt{-1}]$ and $\varepsilon_{1}$ determined by
    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=2-\sqrt{-1},&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=\sqrt{-1},\end{array}$
    in $\mathbb{Q}[\sqrt{-1}]$. For $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$, let us write $D_{\varepsilon}=D_{y}$ if $\varepsilon(\mu)=y$. Then,
    $D_{y}=(y+1-\sqrt{-1}y^{-1})^{2}-4(y-1)=y^{2}-2y-2\sqrt{-1}y^{-1}-y^{-2}+(5-2\sqrt{-1}),$
    and $y$ varies in the set {±$\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$. Noting that $D_{\sqrt{-1}y^{-1}}=D_{y}$, we can compute that $D_{y}$ takes its value in the set
    ${S\coloneqq\left\{\frac{19}{4}-3\sqrt{-1},\frac{51}{4}-\sqrt{-1},\frac{1}{4}-6\sqrt{-1},\frac{9}{4}+2\sqrt{-1}\right\}.}$
    In general, suppose that $D=a+\sqrt{-1}b$ for $a,b\in\mathbb{Q}$ satisfies $D\in\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$. Then, $a^{2}+b^{2}=D\cdot\bar{D}\in\{r^{2}\mid r\in\mathbb{Q}\}$, where $\bar{D}$ is the complex conjugate of $D$. However, for any $D=a+\sqrt{-1}b\in S$, we can check that $a^{2}+b^{2}\notin\{r^{2}\mid r\in\mathbb{Q}\}$. Therefore, $D_{y}\notin\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$ for any $y\in\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$.
  2. 2.
    Choose $\mathbb{k}=\mathbb{F}_{5}\cong\mathbb{Z}/5$ and $\varepsilon_{1}$ determined by
    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1,\end{array}$
    in $\mathbb{F}_{5}$, similar to the one in the proof of Theorem 3.10. In general, if $D\in\mathbb{Z}/5$ satisfies $D\equiv k^{2}\mod 5$ for some $k\in\mathbb{Z}$, then $D\equiv 0,1$ or $4\mod 5$. However, for $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$,
    $D_{\varepsilon}\equiv 33\cdot 4\equiv 2\text{ or }D_{\varepsilon}\equiv 57\cdot 4\equiv 3\mod 5.$
    Therefore, $D_{\varepsilon}\notin\{k^{2}\mid k\in\mathbb{F}_{5}\}$.
Remark 3.12.
To the author’s knowledge, it seems difficult to show Theorem 3.10, even when $K_{0}=3_{1}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$. Indeed, for the trefoil knot $3_{1}$ and the unknot $O$,
$\left.HC_{0}(3_{1})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][Y]/(Y^{2}-\mu Y+\lambda\mu^{3}(1-\mu),\ Y^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-1)),$
@@
$\left.HC_{0}(3_{1})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon Y\mapsto\mu-\mu^{2}.$

### v2

@@ 3.2.3 Application of $\mathbb{Q}$-valued augmentations
For every $m\in\mathbb{Z}_{\geq 1}$, $\deg\bar{h}_{m}=\deg\bar{k}_{m}=m$, and $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ are coprime polynomials in $\mathbb{C}[T]$. In particular, if we define
$r_{m}\coloneqq\frac{T\cdot\bar{k}_{m}}{\bar{h}_{m}}\in\mathbb{C}(T),$
then $\deg r_{m}=\deg(T\cdot\bar{k}_{m})=m+1$ for every $m\in\mathbb{Z}_{\geq 1}$.
@@
By Lemma 3.9, we have $\bar{h}_{0}=1-y_{0}$, $\bar{k}_{0}=1$ and
$\begin{array}[]{cc}\bar{h}_{m+1}=\bar{h}_{m}-T\cdot\bar{k}_{m},&\bar{k}_{m+1}=\bar{h}_{m}+(1-T)\cdot\bar{k}_{m},\end{array}$
for every $m\in\mathbb{Z}_{\geq 0}$. Then, it is easy to see that:
  • For every $m\in\mathbb{Z}_{\geq 1}$, both $\bar{h}_{m}$ and $\bar{k}_{m}$ have $(-1)^{m}\cdot T^{m}$ as the term with the highest degree.
  • For every $m\in\mathbb{Z}_{\geq 0}$, $\bar{h}_{m}=(1-T)\cdot\bar{h}_{m+1}+T\cdot\bar{k}_{m+1}$ and $T\cdot\bar{k}_{m}=T\cdot(\bar{k}_{m+1}-\bar{h}_{m+1})$.
From the first assertion, $\deg\bar{k}_{m}=\deg\bar{h}_{m}=m$ for every $m\in\mathbb{Z}_{\geq 1}$ From the second assertion, if we suppose that $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ have a common root $T=c$ in $\mathbb{C}$ for some $m\in\mathbb{Z}_{\geq 1}$, then $\bar{h}_{m^{\prime}}(c)=c\cdot\bar{k}_{m^{\prime}}(c)=0$ for every $m^{\prime}\in\{0,\dots,m\}$, which is a contradiction when $m^{\prime}=0$ since $1-y_{0}\neq 0$. Thus, $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ are coprime in $\mathbb{C}[T]$ for every $m\in\mathbb{Z}_{\geq 1}$. ∎
Combining Proposition 3.5 and Proposition 3.10 with $\mathbb{k}=\mathbb{Q}$, we obtain the following theorem.
@@
Let $K_{0}=K\#T_{(2,2m+1)}$, where $K$ is an arbitrary knot in $\mathbb{R}^{3}$ and $m\in\mathbb{Z}\setminus\{0,-1\}$. Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
@@
For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. Assuming that $K_{0}=K\#T_{(2,2m+1)}$, we will deduce contradictions in two cases: the case $m\geq 1$ and the case $m\leq-2$.
(The case $m\geq 1$) By Proposition 3.5, for every $Z_{0}\in\{r\in\mathbb{Q}^{*}\mid r^{n}\neq y_{0}\text{ for every }n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=y_{0}$ and $\varepsilon(U)=Z_{0}$. Then, it follows from Proposition 3.10 that $\left.P_{m}\right|_{\mu=y_{0},U=Z_{0}}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$. From the definition (23) of $P_{m}$, it is equivalent to that the equation for $T$
$(Z_{0}-y_{0})\cdot\bar{h}_{m}(T)+y_{0}T\cdot\bar{k}_{m}(T)=0$ (24)
has a solution in $\mathbb{Q}$.
As a polynomial in $\mathbb{Q}(Z)[T]$, let us consider $(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}$. Let $p_{1},\dots,p_{e}\in\mathbb{Q}(Z)[T]$ be irreducible elements such that
$(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}=p_{1}\cdots p_{e}.$
We claim that the degree of $p_{i}$ as a polynomial in $T$ is greater than or equal to $2$ for every $i\in\{1,\dots,e\}$. It is proved by a contradiction as follows: Assume that there exists $i\in\{1,\dots,e\}$ such that $p_{i}=q^{\prime}(T-q)$ for some $q,q^{\prime}\in\mathbb{Q}(Z)$ with $q^{\prime}\neq 0$. Then, we have an equation
$(y_{0}^{-1}Z-1)\cdot\left.\bar{h}_{m}\right|_{T=q}+q\cdot\left.\bar{k}_{m}\right|_{T=q}=0$
in $\mathbb{C}(Z)$. Since $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ does not have a common root in $\mathbb{C}$ by Lemma 3.14, $q$ cannot be a constant and thus, $\left.\bar{h}_{m}\right|_{T=q}\neq 0$ in $\mathbb{C}(Z)$. Therefore, we obtain an equation in $\mathbb{C}(Z)$
$\left.r_{m}\right|_{T=q}=1-y_{0}^{-1}Z$
for $r_{m}=(T\cdot\bar{k}_{m})/\bar{h}_{m}\in\mathbb{C}(T)$. We can compute that
${\deg r_{m}\cdot\deg q=\deg\left(\left.r_{m}\right|_{T=q}\right)=\deg(1-y_{0}^{-1}Z)=1.}$
For the first equality, see Remark 3.13. However, $\deg r_{m}\geq 2$ by Lemma 3.14, and we get a contradiction.
By Hilbert’s irreduciblity theorem (see e.g. [11, Chapter 9, Corollary 2.4], which shows that $\mathbb{Q}$ is a Hilbertian field), there are infinitely many $Z_{0}\in\mathbb{Q}$ such that all of the specializations
$\left.p_{1}\right|_{Z=Z_{0}},\dots,\left.p_{e}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$
are irreducible in $\mathbb{Q}[T]$. We note that the set $\bigcup_{i=1}^{e}\{Z_{0}\in\mathbb{Q}\mid\deg(\left.p_{i}\right|_{Z=Z_{0}})\leq 1\}$ is finite from the above observation of the degree of $p_{1},\dots,p_{m}$ in $T$. In addition, for the fixed rational number $y_{0}\neq 1$, the set $\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ is also finite. As a consequence, there are infinitely many $Z_{0}\in\mathbb{Q}^{*}\setminus\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ such that $\left.p_{i}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$ is irreducible and $\deg(\left.p_{i}\right|_{Z=Z_{0}})\geq 2$ for every $i\in\{1,\dots,e\}$, which means that the equation (24) has no solution in $\mathbb{Q}$. This contradicts Proposition 3.5 and Proposition 3.10.
(The case $m\leq-2$) Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
@@
$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ intersects $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ cleanly along $\overline{K_{1}}$. Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $T_{(2,-2m-1)}$, this contradicts to the result from the above case. ∎
@@
To the author’s knowledge, it seems difficult to show Theorem 3.15, even when $K_{0}=T_{(2,3)}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$. Indeed, for the trefoil knot $T_{(2,3)}$ and the unknot $O$,
$\left.HC_{0}(T_{(2,3)})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(X^{2}-\mu X+\lambda\mu^{3}(1-\mu),\ X^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-1)),$
@@
$\left.HC_{0}(T_{(2,3)})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon X\mapsto\mu-\mu^{2}.$